OpenStudy (anonymous):

divide? (-4q^2 - 22q +12) ÷ (2q+1)

OpenStudy (anonymous):

\[(-4q ^{2} - 22q + 12) \div (2q+1)\]

OpenStudy (anonymous):

take reciprocal remember a/b ÷ c/d = a/b * d/c

OpenStudy (anonymous):

however this is more like a/b ÷ c/1 = a/b * 1/c

OpenStudy (anonymous):

@timo86m how would you find the reciprocal for (2q+1), though?

OpenStudy (anonymous):

oh you just put both terms over 1. @timo86m

OpenStudy (anonymous):

(2q+1) is actually (2q+1)/1 division by one doesn't change it. Thus 1/(2q+1)

OpenStudy (anonymous):

@timo86m okay i get it.. what do i do next?

OpenStudy (anonymous):

write it out :) and we'll see what to do.

OpenStudy (anonymous):

@timo86m okay well